# Gaussian Grouping: Segment and Edit Anything in 3D Scenes

### v1

@@ 4.2 Multiview Reconstruction and Segmentation
Limitation Due to the lack of dynamic modeling and time-dependent updating, the proposed 3D Gaussian Grouping method is currently limited to the static 3D scene. Also, it would also be interested to further explore fully unsupervised 3D Gaussian grouping in the future.
@@ 5 Conclusion
We propose Gaussian Grouping, the first 3D Gaussian-based approach to jointly reconstruct and segment anything in the open-world 3D scene. We introduce an Identity Encoding for 3D Gaussians that is supervised both by 2D mask predictions from SAM and 3D spatial consistency. Based on this grouped and discrete 3D scene representation, we further show it can support versatile scene editing applications, such as 3D object removal, 3D object inpainting and scene recomposition, with both high-quality visual effect and good time efficiency.
@@ 6 Appendix
In this supplementary material, we first conduct additional experiment analysis of our Gaussian Grouping in Section 6.1, including 3D object style transfer and open-world segmentation comparison. Then, in Section 6.2, we describe the detailed process on annotating our proposed LERF-Mask datasets with the visualization of annotation examples. We further provide more detailed 3D object inpainting and style transfer pipeline description in Section 6.3. Finally, we illustrate the algorithm pseudocode of our Gaussian Grouping and more implementation details in Section 6.4, including the method limitation analysis. Please refer to our project page for extensive 3D results comparison.
Figure 10: 3D object style transfer results comparison between Instruct-NeRF2NeRF [10] and our Gaussian Grouping under the same instruction prompt “Turn the bear into a panda” using InstructPix2Pix [2]. Our Gaussian Grouping produces more coherent and natural transfer results across views. Because Gaussian Grouping models the accurate mask of the bear, the spatial details of background region (outside the bear) are more faithfully preserved by our method.
@@
Figure 12: Segmentation comparison between SA3D [4] and our Gaussian Grouping on the rendering view. We adopt PCA to visualize the rendered Identity Encoding features in the rightmost column. Note that SA3D does not support concurrent multi-object segmentation due to its design limitation in the inverse rendering, which requires training and rendering for each segmentation target for around 35 minutes ($\sim$20min for training and $\sim$15min for rendering). In contrast, our Gaussian Grouping shows great efficiency by segmenting all objects of the scene only in 9 minutes.
6.1 Supplementary Experiments
3D Object Style Transfer Comparison Gaussian Grouping supports 3D object style transfer efficiently. We compare with the recent Instruct-NeRF2NeRF [10] on 3D object style transfer in Figure 10, using the same instruction prompt “turn the bear into a panda” and the same image guidance by InstructPix2Pix [2]. Gaussian Grouping produces more coherent and natural transferred “bear” across views. Since our Gaussian Grouping models “anything masks” of open-world 3D scene, including the bear, the spatial details of background regions (outside the bear) are also faithfully preserved by our method. While for Instruct-NeRF2NeRF, the large portion of background regions are unnecessarily getting blurry with degraded quality. We describe the detailed pipeline of 3D object style transfer in Section 6.3.
Segmentation Efficiency Comparison to SA3D We also compare the segmentation results with SA3D [4] on the proposed LERF-Mask dataset for open-world 3D segmentation. Our Gaussian Grouping shows great advantages in segmentation efficiency. Using the same machine, our Gaussian Grouping jointly segments all objects of the 3D scene in 9 minutes, while SA3D requires 35 minutes for each object due to its inverse rendering design in 3D voxel grids. To segmenting each object of the 3D scene, SA3D repeatedly needs a separate new training, which makes SA3D time-consuming and not user-friendly in multi-object segmentation or editing scenarios.
6.2 Details on the LERF-Mask Annotation
Annotation Pipeline To measure the segmentation or fine-grained localization accuracy in open-wold 3D scene, we construct the LERF-Mask dataset based on the existing LERF-Localization [14] evaluation dataset, where we manually annotate three scenes from LERF-Localization with accurate masks instead of using coarse bounding boxes. For each 3D scene, we provide 7.7 text queries with corresponding GT mask label in average. We use Roboflow platform for label annotation, and it uses SAM [15] as an auxiliary tool for interactive segmentation. Similar to the annotation used in LERF, for each of the 3 scenes, we choose 2-4 novel views for testing and annotating the rendering of novel views.
6.3 Steps of Object Inpainting & Style Transfer
3D Object Inpainting Pipeline For inpainting, we remove the 3D Gaussians of the selected target by using a Gaussian Grouping model well-trained for 3D reconstruction and segmentation and add new Gaussians for finetuning. The steps are as follows:
@@ 6.3 Steps of Object Inpainting & Style Transfer
Step 3). On the rendering views after the deletion of the object, we detect the “blurry hole” with Grounding-DINO [24] as the mask for 2D inpainting and use DEVA [7] for association. We use LAMA [41] inpainting on each view as the target for finetuning.
@@
3D Object Style Transfer Pipeline For style transfer, we finetune the 3D Gaussians belong to the correpsonding target by using a Gaussian Grouping model well trained for 3D reconstruction and segmentation. The steps are as follows:
@@
Step 2). Select the target object for style transfer. For each Identity Encoding associated with a 3D Gaussian, we acquire its linear layer classification result. Subsequently, we only finetune those 3D Gaussians that are classified as the label of the selected object. Also, we finetune the 3D Gaussians with position inside the convex hull of the object Gaussians. The Gaussians irrelevant to the editing target are frozen. During finetuning, we freeze the 3D position of Gaussains and make other Gaussian parameters (color, variance, opacity, etc.) trainable.
@@ 6.4 More Implementation Details
More implementation details We implement Gaussian Grouping based on Gaussian Splatting [13]. We add a 16-dimension identity encoding as a feature of each Gaussian, and implement forward and backward cuda rasterization similar to the direct current of Spherical Harmonics. The 3D Identity Encoding has a shape of $N*1*16$, where $N$ is the number of Gaussians. We set the degree of Spherical Harmonics to zero since instance identity does not change across views. The rendered 2D Identity Encoding has a shape of $16*H*W$. 3D Identity Encoding and 2D Identity Encoding share the same identity classification linear layer with $(16,256)$ input and output channels.
During training, we set $\lambda_{\text{2d}}=1.0$ and $\lambda_{\text{3d}}=2.0$. We use the Adam optimizer for both Gaussians and linear layer, with a learning rate of 0.0025 for identity encoding and 0.0005 for linear layer. For 3D regularization loss, we set nearest neighboring number to $k=5$, and sampling points number to $m=1000$. To improve efficiency and avoid calculating loss at boundary points, we downsample the point cloud to 300K to calculate the loss. All datasets are trained for 30K iterations on one A100 GPU.
Algorithm 1 Gaussian Grouping
$w$
, $h$: width and height of the training images
we employ a well-trained zero-shot tracker [7] to propagate and associate masks across views.
we follow the same Gaussians pruning and densification procedure in Gaussian Splatting [13].
$p\leftarrow$ SfM Points $\triangleright$ 3D Positions
$i\leftarrow 0$triangleright Iteration Count
while not converged do
     $L_{\text{id}}\leftarrow\lambda_{2d}\mathcal{L}_{\text{2d}}(\textbf{E}_{\text{id}},\hat{M})+\lambda_{3d}\mathcal{L}_{\text{3d}}(\textbf{e})$triangleright Identity Grouping Loss, see Eq.3 of the paper for $\mathcal{L}_{\text{3d}}$
     if IsRefinementIteration($i$then
[…]
                  RemoveGaussian()
              end if
              if $\nabla_{p}L>\tau_{p}$ then $\triangleright$ Densification
                  if $\|S\|>\tau_{S}$ then $\triangleright$ Over-reconstruction
                       SplitGaussian($p_{j},s_{j},\alpha_{j},c_{j},e_{j}$)
                  elsenormal-$\triangleright$ Under-reconstruction
                       CloneGaussian($p_{j},s_{j},\alpha_{j},c_{j},e_{j}$)
                  end if
              end if
         end for
     end if
     $i\leftarrow i+1$
end while
Algorithm Pseudocode We outline the pseudocode for our Gaussian Grouping in Algorithm 1, where we highlight the introduced core components in both red and bold texts. This idea of our Gaussian Grouping is simple to implement, and straightforward but producing an effective 3D representation to efficiently support versatile downstream scene editing tasks.
Limit Analysis Our Gaussian Grouping segments “anything masks” with the assistance of SAM. But the “anything mask labels” by original SAM [15] have no direct semantic language information. We adopt the Grounding-DINO[24] for open vocabulary segmentation to pick the 2D object, and match our anything masks rendering. When some language prompts are very complicated, the Grounding-DINO can not acquire the correct mask from the input text prompt and will give a wrong mask prediction. In this case, even if we provide the correct mask in anything mask rendering, we do not obtain explicit category information. Also, the zero-shot 2D association accuracy of DEVA [7] will also limit the open-world 3D segmentation performance of Gaussian Grouping. This can be solved by further improvement of the vision language detection model and better association schemes in the future.

### v2

@@ 4.4.5 Limitation
Due to the lack of dynamic modeling and time-dependent updating, the proposed 3D Gaussian Grouping method is currently limited to the static 3D scene. Also, it would also be interesting to further explore fully unsupervised 3D Gaussian grouping in the future.
@@ 5 Conclusion
We propose Gaussian Grouping, the first 3D Gaussian-based approach to jointly reconstruct and segment anything in the open-world 3D scene. We introduce an Identity Encoding for 3D Gaussians that is supervised both by 2D mask predictions from SAM and 3D spatial consistency. Based on this grouped and discrete 3D scene representation, we further show it can support versatile scene editing applications, such as 3D object removal, 3D object inpainting, 3D object style transfer and scene recomposition, with both high-quality visual effects and good time efficiency.
@@ 6 Appendix
In this supplementary material, we first conduct additional experiment analysis of our Gaussian Grouping in Section 6.1, including multi-granularity, segmentation efficiency, quantitative editing evaluation and robustness. Then, in Section 6.2, we describe the detailed process of annotating our proposed LERF-Mask datasets with the visualization of annotation examples. We further provide a more detailed 3D object inpainting and style transfer pipeline description in Section 6.3. Finally, we illustrate the algorithm pseudocode of our Gaussian Grouping and more implementation details in Section 6.4, including the method limitation analysis. Please refer to our project page for extensive 3D results comparison.
6.1 Supplementary Experiments
Segmentation Efficiency In Fig. 13, we compare the segmentation results with SA3D [4] on the proposed LERF-Mask dataset for open-world 3D segmentation. Our Gaussian Grouping shows great advantages in segmentation efficiency. Using the same running environment, our Gaussian Grouping jointly segments all objects of the 3D scene in 9 minutes, while SA3D requires 35 minutes for each object due to its inverse rendering design in 3D voxel grids. To segment each object of the 3D scene, SA3D repeatedly needs separate new training, which makes SA3D time-consuming and not user-friendly in multi-object segmentation or editing scenarios.
Figure 13: Segmentation comparison between SA3D [4] and our Gaussian Grouping on the rendering view. We adopt PCA to visualize the rendered Identity Encoding features in the rightmost column. Note that SA3D does not support concurrent multi-object segmentation due to its design limitation in the inverse rendering, which requires training and rendering for each segmentation target for around 35 minutes ($\sim$20min for training and $\sim$15min for rendering). In contrast, our Gaussian Grouping shows great efficiency by segmenting all objects in the scene only in 9 minutes.
Multi-Granularity of Masks As in Fig. 14, our method can process SAM’s anything masks at different granularity levels in the multiple views as follows:
Step 1). Firstly, SAM predicts dense anything mask proposals (including both large-granularity and small-granularity ones) for each frame.
Step 2). Then, these masks are scored and sorted based on their mask areas. Masks are ranked in descending (from large to small) or ascending (from small to large) order to pick varying levels of granularity for the label of each pixel.
Step 3). Furthermore, we filter large-area or small-area masks through their overlapping IoU with a threshold.
Step 4). For mask association, masks are temporally propagated in a bi-direction to obtain in-clip consensus. Refer to Sec 3.2 of Deva [7] for details.
Figure 14: Consistent tracking results at multi-granularity masks. The 1st row is in coarse granularity, and the 2nd row is fine-grained.
Quantitative Evaluation of Editing Following Instruct-NeRF2NeRF [12], we provide the CLIP Text-Image Direction Similarity evaluation for three of our editing tasks in Table 5. Gaussian grouping supports versatile scene editing tasks, including 3D Object Inpainting, 3D Object Style Transfer and 3D Object Removal. For each of these editing tasks, Gaussian Grouping outperforms the corresponding SOTA method specifically designed for it. For inpainting, the descriptions of original and edited scenes are "Lego toy on the table" and "A flat table". For style transfer, the descriptions are "A bear statue" and "Turn the bear into a panda". For object removal, the descriptions are "A truck on urban street" and "Urban street".
Table 5: Quantitative comparison on three Scene Editing Tasks using the CLIP Text-Image Direction Similarity [12].
Task Dataset Model CLIP Text-Image Direction Similarity
3D Object Inpainting MipNeRF360/Kitchen SPIN-NeRF [34] 0.126
Ours 0.153
3D Object Style Transfer Instruct-NeRF2NeRF/Bear Instruct-NeRF2NeRF [12] 0.171
Ours 0.178
3D Object Removal Tandt/Truck DFFs [19] 0.166
Ours 0.183
Robustness of New Objects SAM+Tracking can process new objects on subsequent frames. When new high-confident segmentation doesn’t match the previous objects and has high confidence, a new instance ID is assigned to it (see the new chair in the red circle in Fig. 15).
Figure 15: Tracking results for a new object in subsequent frames. The red circle highlights the new chair appearing in the 2nd frame.
Sparse View Input For few-shot or sparse-view input, video tracking still obtains a good result for pre-processing, since it is visual feature-based and camera-motion robust. In Figure 16, we use a 3-view input for DEVA pre-processing and still get a decent tracking result. 3D reconstruction of the original Gaussian Splatting is not good with sparse-view input, but it is beyond the scope of this paper. It is retained as a component for sparse-view Gaussian Splatting reconstruction, intended for further research in the future.
Figure 16: Sparse 3-view tracking and reconstruction comparison.
6.2 Details on the LERF-Mask Annotation
Annotation Pipeline To measure the segmentation or fine-grained localization accuracy in the open-wold 3D scene, we construct the LERF-Mask dataset based on the existing LERF-Localization [16] evaluation dataset, where we manually annotate three scenes from LERF-Localization with accurate masks instead of using coarse bounding boxes. For each 3D scene, we provide 7.7 text queries with corresponding GT mask labels on average. We use Roboflow [10] platform for label annotation, and it uses SAM [17] as an auxiliary tool for interactive segmentation. Similar to the annotation used in LERF, for each of the 3 scenes, we choose 2-4 novel views for testing and annotating the rendering of novel views.
@@ 6.2 Details on the LERF-Mask Annotation
6.3 Local Gaussian Editing: Steps of Object Inpainting & Style Transfer
6.3.1 3D Object Inpainting Pipeline
For inpainting, we remove the 3D Gaussians of the selected target by using a Gaussian Grouping model well-trained for 3D reconstruction and segmentation and add new Gaussians for finetuning. The steps are as follows:
@@ 6.3 Local Gaussian Editing: Steps of Object Inpainting & Style Transfer
Step 3). On the rendering views after the deletion of the object, we detect the “blurry hole" with Grounding-DINO [26] as the mask for 2D inpainting and use DEVA [7] for association. We use LAMA [46] inpainting on each view as the target for finetuning.
@@
6.3.2 3D Object Style Transfer Pipeline
For style transfer, we finetune the 3D Gaussians belonging to the corresponding target by using a Gaussian Grouping model well-trained for 3D reconstruction and segmentation. The steps are as follows:
@@
Step 2). Select the target object for style transfer. For each Identity Encoding associated with a 3D Gaussian, we acquire its linear layer classification result. Subsequently, we only finetune those 3D Gaussians that are classified as the label of the selected object. Also, we finetune the 3D Gaussians with position inside the convex hull of the selected object. The Gaussians irrelevant to the editing target are frozen. During finetuning, we freeze the 3D position of Gaussains and make other Gaussian parameters (color, variance, opacity, etc.) trainable.
@@ 6.4 More Implementation Details
6.4.1 More implementation details
We implement Gaussian Grouping based on Gaussian Splatting [15]. We add a 16-dimension identity encoding as a feature of each Gaussian, and implement forward and backward cuda rasterization similar to the direct current of Spherical Harmonics. The 3D Identity Encoding has a shape of $N*1*16$, where $N$ is the number of Gaussians. We set the degree of Spherical Harmonics to zero since instance identity does not change across views. The rendered 2D Identity Encoding has a shape of $16*H*W$. 3D Identity Encoding and 2D Identity Encoding share the same identity classification linear layer with $(16,256)$ input and output channels.
During training, we set $\lambda_{\text{2d}}=1.0$ and $\lambda_{\text{3d}}=2.0$. We use the Adam optimizer for both Gaussians and the linear layer, with a learning rate of 0.0025 for identity encoding and 0.0005 for the linear layer. For 3D regularization loss, we set the nearest neighboring number to $k=5$, and the sampling points number to $m=1000$. To improve efficiency and avoid calculating loss at boundary points, we downsample the point cloud to 300K to calculate the loss. 3D regularization loss only affects the segmentation of identity encoding and does not affect the density of the Gaussians. We use the same adaptive density control as Gaussian Splatting. All datasets are trained for 30K iterations on one A100 GPU.
6.5 Limitation
Our Gaussian Grouping segments “anything masks” with the assistance of SAM. But the “anything mask labels" by original SAM [17] have no direct semantic language information. We adopt the Grounding-DINO[26] for open vocabulary segmentation to pick the 2D object, and match our anything masks rendering. When some language prompts are very complicated, the Grounding-DINO can not acquire the correct mask from the input text prompt and will give a wrong mask prediction. In this case, even if we provide the correct mask in anything mask rendering, we do not obtain explicit category information. Also, the zero-shot 2D association accuracy of DEVA [7] will also limit the open-world 3D segmentation performance of Gaussian Grouping. This can be solved by further improvement of the vision language detection model and better association schemes in the future.